# Where’s the Bug? Attention Probing for Scalable Fault Localization

### v1

@@ 1 Introduction
Automated software fault localization (FL) aims to help a programmer answer the question, “Where’s the bug?”, ideally pointing to specific lines of buggy code. The traditional FL approaches rely on executable tests to determine the buggy lines (Johnson et al., 2013). Without relying on tests, FL is even more challenging since such a system must reason about what is buggy without external feedback. But recently, supervised training of models on large, labeled datasets (Yang et al., 2024; Meng et al., 2022), and prompting of the largest LLMs such as GPT-4o has shown promise at FL without tests (Wu et al., 2023). On a single method context, large-scale supervised training and prompting approaches can significantly surpass the traditional techniques which need tests (Yang et al., 2024; Wu et al., 2023).
@@
  • We propose Bug Attention Probe (BAP) as a general method for scalable fault localization, requiring only coarse-grained detection supervision and eschewing the need for localization labels (Section LABEL:sec:ourmethod).
@@ 2.2 LLM Probing
These datasets, $\mathcal{H}_{\text{Det}}$ and $\mathcal{H}_{\text{Loc}}$, are not directly amenable to standard probing since each sample, $\text{LLMRep}(p,k)$, is a sequence of hidden representations varying in length across samples. Producing a general-purpose fixed-length sequence representation from an autoregressive LLM is a challenging problem (Liu et al., 2023), so probing methods typically apply a simple pooling operation, $\text{POOL}:\mathbb{R}^{T\times d}\rightarrow\mathbb{R}^{1\times d}$, such as selecting the last token or averaging all tokens. Specific to using $\mathcal{H}_{\text{Loc}}$, existing work takes the approach of converting the labels $l(p)$ into a binary mask over the buggy program lines so that a model can be trained to predict such masks (Yang et al., 2024).
@@ 2.4 RQ1: FL Performance
FL performance of BAP and baselines on eight datasets is shown in Table LABEL:tab:defects4j. To compare with existing work on Defects4J, we perform 10-fold cross validation using method-level context following (Wu et al., 2023). Results for DeepFL, TRANSFER-FL, and Smart-FL are from Yang et al. (2024) and Zeng et al. (2022). Note that “n/a” results indicate baselines which could not be evaluated due to requiring tests which were only available for Defects4J.
@@ 3 Related Work
Automated Fault Localization. Methods for FL include the traditional spectrum-based (SBFL) and mutation-based (MBFL) methods which require executable code and deep-learning based approaches (Wong et al., 2016b). SBFL methods are simple but have low accuracy while MBFL and deep learning approaches have higher accuracy at larger computational cost (Wong et al., 2016b). Various deep learning approaches combine SBFL and MBFL with semantic features from deep models (Li et al., 2019; Meng et al., 2022; Zeng et al., 2022). Recently, LLMs have significantly outperformed SBFL and MBFL approaches on FL on the method level (Wu et al., 2023). Prompting and agent-based systems can even perform repository-level FL (Xia et al., 2024; Qin et al., 2024), but they must reduce the problem to method-level FL (Xia et al., 2024). LLMAO (Yang et al., 2024) trains an adapter on an LLM from strong FL supervision to perform FL without executable tests, and WELL (Zhang et al., 2024) finetunes an LLM on bug detection supervision and interprets the attention for FL. Unlike these approaches, our method uses LLM probing, and we leverage bug detection supervision to scale to more available data.
@@ A.1 Datasets
The detailed breakdown of each of the datasets we used, other than Defects4J. We split datasets into groups based on the three domains of syntax, single line bugs, and vulnerabilities and show the breakdown in Table 4, 5, and 6 respectively.
@@ B.1 Error Bars for FL Results
Error bars for the results in Table LABEL:tab:defects4j are provided in Table 7. The prompting methods have no error bars because we use greedy decoding.

### v2

@@ 1 Introduction
Automated software fault localization (FL) aims to help a programmer answer the question, “Where’s the bug?”, ideally pointing to specific lines of buggy code. The traditional FL approaches rely on executable tests to determine the buggy lines (Johnson et al., 2013). Without relying on tests, FL is even more challenging since such a system must reason about what is buggy without external feedback. But recently, supervised training of models on large, labeled datasets (Yang et al., 2024a; Meng et al., 2022), and prompting of the largest LLMs such as GPT-4o has shown promise at FL without tests (Wu et al., 2023). On a single method context, large-scale supervised training and prompting approaches can significantly surpass the traditional techniques which need tests (Yang et al., 2024a; Wu et al., 2023).
@@
  • We propose Bug Attention Probe (BAP) as a general method for scalable fault localization, requiring only coarse-grained detection supervision and eschewing the need for localization labels (Section 3).
@@ 2.2 LLM Probing
These datasets, $\mathcal{H}_{\text{Det}}$ and $\mathcal{H}_{\text{Loc}}$, are not directly amenable to standard probing since each sample, $\text{LLMRep}(p,k)$, is a sequence of hidden representations varying in length across samples. Producing a general-purpose fixed-length sequence representation from an autoregressive LLM is a challenging problem (Liu et al., 2023), so probing methods typically apply a simple pooling operation, $\text{POOL}:\mathbb{R}^{T\times d}\rightarrow\mathbb{R}^{1\times d}$, such as selecting the last token or averaging all tokens. Specific to using $\mathcal{H}_{\text{Loc}}$, existing work takes the approach of converting the labels $l(p)$ into a binary mask over the buggy program lines so that a model can be trained to predict such masks (Yang et al., 2024a).
@@ 2.3 Challenges
3 Attention Probing for Fault Localization
In this section, we introduce our approach for addressing the previous challenges.
We propose the Bug Attention Probe (BAP), an LLM probing technique for performing FL which provides scalability through lightweight requirements (both in terms of training supervision and model size), interpretable code-level localization, and handling of multi-line bugs. We take inspiration from attention probing (Tenney et al., 2018; Niu et al., 2022), a technique from the interpretability literature for studying linguistic phenomena such as the attention placed on verb tokens, but we use such an approach for code-level FL, as in localizing bugs to lines within code.
We illustrate BAP, its process of training from bug detection data, $\mathcal{H}_{\text{Det}}$, and the computation of line-level bug localization from attention weights in Figure 2.
3.1 Why Attention Probing for Fault Localization?
Our main motivation for using an attention mechanism is that the attention pooling operation trained for the task of bug detection encourages the probe to learn to attend to the location of bugs, without using strong FL supervision. Intuitively, the probe attends to parts of the code which are more informative for bug detection, and we hypothesize that these locations often correspond to the location of the bugs.
For bugs which are localized to multiple lines, which is the majority of real world faults, an attention mechanism is also helpful since it operates on the token-level, making no distinction between a single and multi-line bug location.
Finally, an attention mechanism is efficient since it operates over all tokens in parallel. This means that a complete ranking of all the program lines in terms of the likelihood they contain a bug is produced by a single forward pass of attention. This is in contrast to methods such as LLM prompting which must continue to output more tokens to localize a bug to more lines.
3.2 From Weak Supervision to Token-level Localization
To enable learning FL from weak supervision, we use a single layer Transformer decoder block as the architecture of BAP to factorize the bug detection task into localization (attention) over tokens and detection (classification) on the sequence-level. The input to the probe consists of a sequence of token representations from the $k$th layer of the LLM, $\text{LLMRep}(p,k)=[z_{1},\dots,z_{T}]\in\mathbb{R}^{T\times d}$, for a program $p$ consisting of $T$ tokens. The standard multi-head attention mechanism takes $\text{LLMRep}(p,k)$ as input and outputs attention scores $a\in\mathbb{R}^{M\times T\times T}$ where $M$ is the number of attention heads, and processed token representations $v\in\mathbb{R}^{T\times d}$. We compute token-level attention scores of all tokens to the final token as $\bar{a}=\frac{1}{M}\sum_{m=1}^{M}a_{m,-1,:}\in\mathbb{R}^{T}$ where we average the attention from each head. The process of producing token-level attention from input code is also shown in lines 1-5 of Algorithm 1.
To get useful token-level attention from our model, we must train it on a downstream task. To train BAP, we use weak supervision in the form of a bug detection dataset, $\mathcal{H}_{\text{Det}}$ defined in Section 2. We pass the last token from the processed sequence $v$ from the attention through a feed-forward network (MLP) to produce a single scalar output representing the buggy/non-buggy prediction. The model is then optimized using gradient descent with the binary cross-entropy loss. After training the probe for bug detection, we then can examine the learned attention weights $\bar{a}$ which provide a token-level fault localization.
3.3 From Token-level to Code-level Localization
A major challenge with this approach is that tokens are not interpretable for programmers. Therefore, we provide a method to aggregate token-level fault localization into a code-level localization. Our approach is summarized in lines 6 to 8 of Algorithm 1. In our experiments, we focus on line-level granularity, but this method also allows us to perform statement-level and function-level localization without any significant modifications.
Algorithm 1 BAP Line-level Fault Localization
1:  Input: code sample $p$, layer $k$.
2:  Tokenize $p$ to get $p=[t_{1},\dots,t_{T}]$.
3:  $z=\text{LLMRep}(p,k)$
4:  $v,a=\text{MHA}(z)$ where $v\in\mathbb{R}^{T\times d}$ and $a\in\mathbb{R}^{M\times T\times T}$
5:  $\bar{a}=\frac{1}{M}\sum_{m=1}^{M}a_{m,-1,:}$Average attention over all heads, for the last token}
6:  Group attention scores into lines: $s_{1}=[\bar{a}_{1},\dots,\bar{a}_{i}]$, $s_{2}=[\bar{a}_{i+1},\dots,\bar{a}_{i+j}],\dots$
7:  $l_{i}=\sum_{t\in s_{i}}t$ for all $i$Line-level attention score}
8:  Return: $\operatorname{argsort}_{i\in[1,L]}l_{i}$
Line 5 of Algorithm 1 computes the token-level attention scores, and we call $\bar{a}_{i}$ the probe’s attention score for the $i$th token. To produce a line-level attention score, we sum the attention scores for all tokens in the $i$th line, $s_{i}=[\bar{a}_{i_{1}},\bar{a}_{i_{2}},\dots]$, to produce $l_{i}$, the probe’s attention score for the $i$th line. The computation of line-level attention scores is shown in line 7 of Algorithm 1.
Line 8 of Algorithm 1 shows that the resulting line-level localization from BAP is the ranking of input lines based on their respective attention scores. In practice, we truncate to the top-$k$ lines out of total lines $L$. The line with the highest attention score is thus noted as the top-1 prediction.
4 Experiments
We evaluate BAP over a diverse suite of eight fault localization benchmarks. This includes Defects4J (Just et al., 2014), the most popular fault localization benchmark, as well as seven additional benchmarks covering three general bug types. In the rest of this section, we first introduce the datasets, then the baseline methods, and finally the results of each experiment. We answer the following research questions in this section:
RQ1
How effective is BAP at FL in diverse scenarios?
RQ2
How does the efficiency of BAP compare to baselines?
RQ3
Can BAP effectively localize multi-line bugs?
RQ4
How does the generalization ability of BAP compare to zero-shot prompting of the base model?
4.1 Datasets
We evaluate on eight datasets summarized below.
Table 1: A summary of the datasets we use for our evaluation. For a further breakdown of buggy samples into categories, see our discussion in Appendix A.1.
# Train # Test
Dataset Bug Clean Bug Clean $\mathbb{E}$[LoC]
Defects4J v1.2.0 368 368 90 90 35.8
Defects4J v3.0.1 n/a n/a 437 n/a 46.7
GitHub-Py 1323 1323 400 400 9.3
GitHub-J 1370 1370 460 460 19.3
DeepFix 1475 1475 365 365 26.3
TSSB 4085 3745 1104 1080 24.8
ManySStuBs4J 3821 3821 1093 1093 15.5
Juliet-J 4039 3061 1011 989 62.5
Juliet-C 3718 3697 966 939 44.7
Table 2: Comparison of BAP with existing FL methods across eight benchmarks. We evaluate line-level localization performance on a method-level context, measured by top-1 accuracy. From left-to-right: Defects4J v1.2.0, GitHub-Python, GitHub-Java, DeepFix, TSSB-3M, ManySStuBs4J, Juliet-Java, and Juliet-C. Error bars are provided in Appendix B.
Method D4J GH-Py GH-J DeepFix TSSB MS4J Juliet-J Juliet-C Avg.
Random 0.144 0.100 0.134 0.038 0.069 0.124 0.025 0.058 0.087
DeepFL 0.144 n/a n/a n/a n/a n/a n/a n/a 0.144
SmartFL 0.158 n/a n/a n/a n/a n/a n/a n/a 0.158
TRANSFER-FL 0.218 n/a n/a n/a n/a n/a n/a n/a 0.218
CodeLlama-70B 0.212 0.145 0.316 0.084 0.077 0.169 0.038 0.095 0.142
Llama-3.3-70B 0.269 0.225 0.272 0.092 0.114 0.211 0.072 0.040 0.162
Qwen2.5-72B 0.157 0.333 0.289 0.124 0.088 0.194 0.061 0.040 0.161
DeepSeek-R1-Distill-Llama-70B 0.221 0.188 0.218 0.035 0.138 0.185 0.041 0.025 0.131
GPT-4o 0.249 0.375 0.365 0.097 0.089 0.240 0.009 0.026 0.181
Linear Probe Llama-3.2-11B 0.279 0.373 0.300 0.140 0.202 0.235 0.048 0.043 0.202
LLMAO-Llama-3.2-11B 0.144 0.190 0.188 0.078 0.118 0.116 0.063 0.113 0.126
WELL-CodeBERT 0.090 0.575 0.532 0.129 0.094 0.111 0.216 0.059 0.226
WELL-Llama-3.2-11B 0.236 0.028 0.139 0.000 0.054 0.081 0.000 0.000 0.067
GridLoc-Llama-3.2-11B 0.291 0.498 0.206 0.332 0.262 0.339 0.158 0.039 0.266
BAP-Llama-3.2-11B 0.334 0.575 0.568 0.481 0.237 0.291 0.096 0.217 0.350
Defects4J Bugs along with commits to fix the bug. We use two versions of the dataset: the standard Defects4J v1.2.0 (Just et al., 2014) dataset containing 395 bugs, and the additional 543 bugs from Defects4J v3.0.1 released in November 2024 which we use as a stronger evaluation of generalization. In both versions, we split each buggy file into a set of buggy methods and evaluate on the method level following Wu et al. (2023).
GithHub-Python (Yasunaga & Liang, 2021) and GithHub-Java (Santos et al., 2018) consist of code mined from GitHub with syntax errors in Python and Java respectively.
DeepFix Real C programs written by students, some containing beginner syntax mistakes (Gupta et al., 2017).
TSSB-3M “Simple, stupid bugs” (SStuBs) in Python which are mined from GitHub and categorized (Richter & Wehrheim, 2022). Despite their name, these bugs are extremely challenging for humans to localize.
ManySStuBs4J Java SStuBs (Karampatsis & Sutton, 2020).
Juliet-Java (jul, 2023b) and Juliet-C (jul, 2023a) consist of synthetic code corresponding to Common Weakness Enumerations (CWEs). We rename variables and function names to remove indicators of the vulnerability. We also remove comments and rename imports that refer to the dataset names. In total, we consider 89 CWEs from both the Juliet datasets.
The datasets are summarized in Table 1 and we provide a detailed breakdown of the datasets in Appendix A.1.
4.2 Baselines
We group the baselines into three types: traditional FL methods that require code execution, LLM prompting of different models, and LLM probing/training.
Traditional FL Methods For methods requiring code execution, we compare with DeepFL (Li et al., 2019), SmartFL (Zeng et al., 2022), and TRANSFER-FL (Meng et al., 2022) on Defects4J since these are the best performing traditional FL methods. Results for DeepFL and Transfer-FL are from Yang et al. (2024a), and we cite SmartFL results directly from Zeng et al. (2022). We can not evaluate these benchmarks on the other datasets since they do not provide tests.
Prompting Methods For models, we use a diverse set of four open-weights LLMs of size $\sim$70B parameters as well as a proprietary model. We consider Llama 3.3 70B (Dubey et al., 2024), Qwen 2.5 72B (Yang et al., 2024b), CodeLlama 70B (Roziere et al., 2023), and DeepSeek-R1-Distill-Llama-70B (Guo et al., 2025) as the main open-weights LLMs and we use GPT-4o (Hurst et al., 2024) as the representative proprietary LLM for prompting experiments. Llama 3.3 70B and Qwen 2.5 72B are LLMs pretrained on a diverse dataset of natural language and code while CodeLlama 70B additionally trained on primarily code data (Roziere et al., 2023). DeepSeek-R1-Distill-Llama-70B is a “reasoning” model in that it can use longer chain-of-thought output to solve more complex problems (Guo et al., 2025).
We experimented with several prompts based on that used by Wu et al. (2023). Our prompt asks for the buggy line text as well as the line number in case the LLM cannot count lines. We use this prompt with temperature 0 sampling for all models. The prompt is provided in Appendix A.2.
Probing Methods Finally, for LLM probing, we consider the following baselines:
  • Linear Probing: performing logistic regression on the last token representation from each code sample to predict if the code is buggy or not. Following Zou et al. (2023), we apply the trained classifier to each token in the sequence for getting token-level scores and then we derive line rankings the same way as BAP.
  • GridLoc (Niu et al., 2022): the only other attention probing method we are aware of, which uses an RNN to learn attention weights. We apply this method to learning the bug detection task from the dataset $\mathcal{H}_{\text{Det}}$, and interpret the resulting attention weights for FL using our line-aggregation method.
  • LLMAO (Yang et al., 2024a): trains an adapter using strong FL supervision. This method was designed for the CodeGen (Nijkamp et al., 2023) models, and we used the largest available CodeGen model. We additionally adapted their code for the Llama-3 family of models and report results on Llama-3.2-11B.
  • WELL (Zhang et al., 2024): the first method to consider learning FL without strong supervision. This method finetunes a bidirectional LLM (CodeBERT (Feng et al., 2020)) for bug detection and then interprets the attention weights of the last layer for fault localization. We additionally report results on Llama-3.2-11B, even though it has exclusively causal attention.
@@ 4.3 RQ1: FL Performance
FL performance of BAP and baselines on eight datasets is shown in Table 2. To compare with existing work on Defects4J, we perform 10-fold cross validation using method-level context following (Wu et al., 2023). Results for DeepFL, TRANSFER-FL, and Smart-FL are from Yang et al. (2024a) and Zeng et al. (2022). Note that “n/a” results indicate baselines which could not be evaluated due to requiring tests which were only available for Defects4J.
@@ 5 Related Work
Automated Fault Localization. Methods for FL include the traditional spectrum-based (SBFL) and mutation-based (MBFL) methods which require executable code and deep-learning based approaches (Wong et al., 2016b). SBFL methods are simple but have low accuracy while MBFL and deep learning approaches have higher accuracy at larger computational cost (Wong et al., 2016b). Various deep learning approaches combine SBFL and MBFL with semantic features from deep models (Li et al., 2019; Meng et al., 2022; Zeng et al., 2022). Recently, LLMs have significantly outperformed SBFL and MBFL approaches on FL on the method level (Wu et al., 2023). Prompting and agent-based systems can even perform repository-level FL (Xia et al., 2024; Qin et al., 2024), but they must reduce the problem to method-level FL (Xia et al., 2024). LLMAO (Yang et al., 2024a) trains an adapter on an LLM from strong FL supervision to perform FL without executable tests, and WELL (Zhang et al., 2024) finetunes an LLM on bug detection supervision and interprets the attention for FL. Unlike these approaches, our method uses LLM probing, and we leverage bug detection supervision to scale to more available data.
@@ A.1 Datasets
The detailed breakdown of each of the datasets we used, other than Defects4J. We split datasets into groups based on the three domains of syntax, single line bugs, and vulnerabilities and show the breakdown in Table 6, 7, and 8 respectively.
@@ B.1 Error Bars for FL Results
Error bars for the results in Table 2 are provided in Table 9. The prompting methods have no error bars because we use greedy decoding.